OpenStudy (anonymous):

What is the length of the base of a triangle with height 11 cm and area 137.5 cm2?

OpenStudy (justuu):

let b=base length 137.5=0.5*11b b=25

OpenStudy (anonymous):

\[137.5 \times 2 \div 11\] so it should be like 250

OpenStudy (anonymous):

Thank you! Now I understand

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so there are two methods ...